George Papageorgiou

PERSONAL INFORMATION

Date of Birth:  31 May 1972
Nationality:  Greek/British (dual nationality)
Gender:  Male
Marital Status:  Single
Work Address:  University of Cambridge, Department of Engineering,

Trumpington Street, Cambridge CB2 1PZ, ENGLAND. Tel: +44-1223-339706 Fax: +44-1223-332662

          E-mail: gp202@eng.cam.ac.uk

Web site: http://www-control.eng.cam.ac.uk/gp202/

EDUCATION

Oct. 1994 - Jul. 1998 Department of Engineering, University of Cambridge, England and Darwin College, Cambridge, England. Degree Awarded: PhD under the supervision of Professor K. Glover. Thesis title: “Robust Control System Design: H_{∞} Loop Shaping and Aerospace Applications.”

Oct. 1991 - Jun. 1994 Faculty of Engineering, University of Bristol, England.

Degree Awarded: Bachelor of Engineering with First Class Honours in Aerospace Engineering.

PRIZES AND AWARDS

1995 - 1998 Defence Evaluation and Research Agency (DERA), Bedford Cooperative Awards in Science and Engineering (CASE) Grant, for post-graduate study carried out at the University of Cambridge, England.

1994 - 1997 Engineering and Physical Sciences Research Council (EPSRC) Studentship, for post-graduate study carried out at the University of Cambridge, England.

1994 Royal Aeronautical Society Prize; the Royal Aeronautical Society Prize is awarded to the top student in the Aerospace Engineering Department, University of Bristol.

1994 Second place in the 1994 Bristol Branch of the Royal Aeronautical Society Lecture Competition; the paper presented was titled “Actively Controlled Two Degrees of Freedom Wind Tunnel Model.”

1993 Collar Prize; the Collar Prize is awarded to the second year aerospace student at the University of Bristol who has combined academic excellence with other activities.

1992 A one-year Undergraduate Scholarship in the Faculty of Engineering, University of Bristol for the session 1993/94.

PUBLICATIONS

Book Chapters:

[BC1] G. Papageorgiou, K. Glover, A. Smerlas and I. Postlethwaite, “H_{∞} Loop Shaping.” Robust Flight Control: A Design Challenge GARTEUR, Volume 224 of Lecture Notes in Control and Information Sciences, Chapter 7, Springer-Verlag, 1997.

[BC2] G. Papageorgiou, K. Glover and R.A. Hyde, “The H_{∞} Loop Shaping Approach.” Robust Flight Control: A Design Challenge GARTEUR, Volume 224 of Lecture Notes in Control and Information Sciences, Chapter 29, Springer-Verlag, 1997.

Refereed Journal Papers:

[JP1] G. Papageorgiou and K. Glover, “Design of a robust gain scheduled controller for the High Incidence Research Model.” Accepted for publication in the AIAA Journal of Guidance, Control, and Dynamics, July 1999.

Refereed Conference Papers:

[CP1] G. Papageorgiou, M. Huzmezan, K. Glover and J. Maciejowski, “Combined MBPC/H_{∞} Autopilot for a Civil Aircraft.” Presented at the 1997 American Control Conference.

[CP2] G. Papageorgiou and K. Glover, “Design and Development of a Wind Tunnel Model for High Angle of Attack Active Control Experiments.” Presented at the 1997 European Control Conference.

[CP3] G. Papageorgiou and K. Glover, “A Systematic Procedure for Designing Non-Diagonal Weights to Facilitate H_{∞} Loop Shaping.” Presented at the 1997 Conference on Decision and Control.

[CP4] G. Papageorgiou and K. Glover, “H_{∞} loop shaping: Why is it a sensible procedure for designing robust flight controllers?” Presented at the 1999 AIAA Guidance, Navigation and Control Conference.

[CP5] G. Papageorgiou and K. Glover, “Design of a robust two degrees-of-freedom controller for an actively controlled wind tunnel model.” Presented at the 1999 AIAA Guidance, Navigation and Control Conference.

[CP6] G. Papageorgiou and K. Glover, “Design of a robust gain scheduled controller for the High Incidence Research Model.” Presented at the 1999 AIAA Guidance, Navigation and Control Conference.

[CP7] G. Papageorgiou and K. Glover, “Design, development and control of the HIRM wind tunnel model.” To be presented at the 1999 Conference on Decision and Control.

Papers in Preparation:

[PP1] G. Papageorgiou and K. Glover, “Design, Development and Testing of a Wind Tunnel Model.” To be submitted to IEEE Transactions on Control Systems Technology, 1999.

[PP2] G. Papageorgiou and K. Glover, “Design of Non-Diagonal Weights to Facilitate H_{∞} Loop Shaping.” To be submitted to Automatica, 1999.

WORK EXPERIENCE

1998 - Research Associate, Department of Engineering, University of Cambridge. Project title: “Design and analysis of robust flight control laws.” The project is sponsored by DERA, Bedford and supervised by Professor K. Glover.

1998 - Laboratory Marker, Department of Engineering, University of Cambridge. Flight Control Experiment (3^{rd} year lab).

1998 Author of Assessed Course Work, Department of Engineering, University of Cambridge. Multivariable and Optimal Control (4^{th} year course).

1996 - Fourth Year Project Supervisor, Department of Engineering, University of Cambridge.

1995 - 1996 Undergraduate Supervisor, Department of Engineering, University of Cambridge. Control and Signal Processing (3^{rd} year course).

1995 - 1998 Laboratory Demonstrator, Department of Engineering, University of Cambridge. Flight Control Experiment (3^{rd} year lab).

1990 -  Seasonal invigilation of exams for the British Council in Athens.

1989 and 1990 Summer vacation job in a hotel in Loutraki, Greece as a receptionist and assistant accountant.

POSITIONS of RESPONSIBILITY

1996 - 1998 Captain of the Darwin College Basketball Club; the Darwin College Basketball Club has about 30 members, and one of the most successful basketball teams in Cambridge.

1995 - 1996 Secretary of the Darwin College Student Association; the Darwin College Student Association is responsible for representing the interests of the students in the college, and coordinating the social life of the students.

1994 - 1995 President of the Cambridge University Hellenic Society; the Cambridge University Hellenic Society has about 150 members, and its aim is to get people in touch with the Hellenic way of life and to promote the Hellenic culture and spirit.

TECHNICAL SKILLS

• Fluent in Greek, English and French.

• Working knowledge of the programming languages Matlab, C and Maple.

• Working knowledge of Latex and Word for Windows.

• Experience with dSPACE GmbH software and hardware (tools for rapid control prototyping and hardware-in-the-loop simulation).

• Experience with UNIX, Windows and DOS operating systems.

PROFESSIONAL MEMBERSHIPS

• Member of the Institute of Electrical and Electronic Engineers (IEEE).

• Student Member of the American Institute of Aeronautics and Astronautics (AIAA).

OTHER INTERESTS

Football, Basketball, Squash, Wind-surfing, Tennis, Table Tennis, Travelling, Reading, Bridge, Playing the synthesiser.

REFEREES

Professor K. Glover,  Dr. R.A. Hyde,
University of Cambridge,  Cambridge Control Ltd,
Department of Engineering,  Cowley Road,
Cambridge CB2 1PZ,  Cambridge CB4 4WZ,
ENGLAND.  ENGLAND.
Tel: +44-1223-332752  Tel: +44-1223-423200/205
Fax: +44-1223-332662  Fax: +44-1223-423289/255

E-mail: kg@eng.cam.ac.uk  E-mail: Rick Hyde@camcontrol.co.uk